# HAF-RM: A Hybrid Alignment Framework for Reward Model Training

### v3

@@ 1 Introduction
Figure 1: The standard reward model substitutes the policy layer from the policy model, while our HaF model retains the policy layer. By optimizing the model’s two outputs, we achieve a better alignment process for the reward model with little additional training overhead.
A reward model is typically structured with two components: a transformer-based model (referred to as the internal preference model), and a projection module called “reward layer” (usually a linear layer with normalization). The former outputs preference vectors for each token, while the latter maps these vectors to sequence-level rewards. The standard practice for training the reward model involves utilizing the ranking loss of paired rewards. However, optimizing both two components using such a single sequence-level objective may cause insufficient supervision for token-level preference modeling. We argue that hybrid optimization of the two components of the reward model with corresponding token-level and sequence-level objectives will lead to more consistent improvements.
@@ 2.2 Basic Loss Functions
where $\sigma\left({\cdot}\right)$ is the sigmoid function (derivation can be found in C.1).
@@ 3.2 Compared Models
First, we compare our framework with the standard training approach, wherein the reward model only has a reward layer dedicated to reward prediction and is optimized only with reward loss, as delineated in Eq. 2.
@@
Although DPO loss (as in Eq. 4) is typically used for training policy models rather than reward models, it can implicitly convert model’s outputs into reward values (Rafailov et al., 2023). Therefore, the DPO-trained model can also function as a reward model (Rafailov et al., 2024). Following the work of Lambert et al. (2024), we also evaluate the model trained with DPO loss.
@@ 4.2.2 RLHF
As shown in Table 5, the improvement of HaF is particularly evident on the Chat dataset, with its win rate increasing throughout the training, highlighting the superiority of the HaF reward model. In contrast, during safety training, the HaF reward model shows a significant advantage over the baseline model primarily in the middle stages of training. This is likely because both models have largely achieved harmless responses on the test set, resulting in minimal differentiation between the two reward models.
@@ 5 Related Work
Reward model was proposed to modeling human language preferences (model that outputs preference values based on questions and answers) (Christiano et al., 2017), then the explosive growth of research on reward models (McKinney et al., 2023) and large language models (Wei et al., 2022; Park et al., 2023; Zheng et al., 2023) emerged after the popularity of ChatGPT.
@@
Although many data-oriented methods have greatly enhanced the performance of reward models, the field of reward model optimization has been rarely explored. Currently, the training of reward models basically follows the process proposed by OpenAI (Christiano et al., 2017). It involves initializing the reward model using a finetuned model, then transforming the model’s predictions into probability values through the Bradley-Terry model, and optimizing these probabilities using cross-entropy loss. Considering the widespread practical applications of reward models, the attention given to their training paradigms does not match their importance.
@@ Appendix A Experiments Setup
lora rank 64 optimizer AdamW precision fp16
@@
To train the reward model, we use DPO Loss as the policy loss in HAF and set policy ratio $\alpha=0.2$. The learning rate is $1.0\times 10^{-5}$ for Phi-2 and Mistral-Instruct with the baseline method, $3.0\times 10^{-5}$ for Mistral-Base and Mistral-Instruct using other methods, and the batch size is set at 16. These configurations are the optimal combination of learning rates (among $1.0\times 10^{-4}$, $3.0\times 10^{-5}$, $1.0\times 10^{-5}$, $3.0\times 10^{-6}$) and batch sizes (among $4$, $16$, $128$). A single RTX A6000 with 48GB memory is used for training the reward model. The model used for testing is the checkpoint that achieves the highest reward on the validation set.
For PPO training in Section 4.2.2, we set the total batch size at 16. The maximum number of new tokens generated is set to 256, and the learning rate is $2.0\times 10^{-5}$. The training is conducted over a maximum of 100,000 episodes. All other settings follow the implementation in the LLaMA-Factory library. The generation config includes top_p=$0.9$, do_sample=True.
@@ Appendix F GPT Judgement
Table 8 shows the consumption approximation for getting top-1, top-2 responses and the complete order out of 4/8 responses. We consider that performing a single sorting operation on eight responses with the model may result in a loss of precision. Besides, while binary comparisons exhibit high accuracy, repeated binary comparisons inevitably lead to cumulative errors and erroneous outcomes. Therefore, whether from a cost or accuracy standpoint, it is not a favorable option. In practice, we obtain the top 2 responses by ranking 4 responses with gpt-4-turbo-2024-04-09 at once. For 8 candidate responses, we first evenly divide them into two groups and use GPT to rank the responses of each group, then we rank the two sets of the top 2 responses to get the top 2 responses among 8 candidates.
@@
Phi-2HAF 28.67 52.51 32.49 53.06 37.46 65.94 45.44 74.25 24.79 45.67
Phi-2DPO 20.85 45.62 17.68 39.71 31.89 56.07 26.42 53.07 17.87 36.67
Phi-2baseline 15.45 34.63 32.85 50.90 27.84 51.64 43.62 73.91 17.41 37.48
MistralHAF 22.86 41.95 14.61 25.38 24.12 42.69 16.39 35.24 12.99 29.23
MistralDPO 14.57 32.91 10.00 24.61 13.68 30.62 13.31 31.14 12.52 25.98
Mistralbaseline 14.82 30.40 13.07 24.61 25.05 43.85 14.95 31.14 12.29 23.20

### v4

@@ 1 Introduction
Figure 1: HaF model structure. It retains the policy layer which outputs the token-level probability.
@@
A reward model is typically structured with two components: a transformer-based model (referred to as the “internal preference model”), and a projection module called “reward layer” (usually a linear layer). The former outputs preference vectors for each token, while the latter maps these vectors to sequence-level rewards. We argue that the standard practice for training the reward model may cause insufficient supervision for preference modeling, which can be improved by performing hybrid supervision of both token-level and sequence-level.
@@ 2.2 Basic Loss Functions
where $\sigma\left({\cdot}\right)$ is the sigmoid function (derivation can be found in Appendix C.1).
@@
Figure 2: HaF training framework. We add the reward layer to the language model while retaining its policy layer. During training, we optimize both the token-level rewards and sequence-level rewards for the input triplets by maximizing the reward differences between better responses and worse responses.
@@ 3.2 Compared Models
We compare our framework with the standard training approach, wherein the reward model only has a reward layer dedicated to reward prediction and is optimized only with reward loss, as delineated in Eq. 2.
@@
DPO can implicitly convert model’s outputs into reward values (Rafailov et al., 2023), so the model can also function as a reward model (Rafailov et al., 2024). Following the work of Lambert et al. (2024), we evaluate the model trained with DPO loss.
@@ 4.2.2 RLHF
As shown in Figure 6, the improvement of HaF is particularly evident on the Chat dataset, with its win rate increasing throughout the training, highlighting the superiority of the HaF reward model. In contrast, during safety training, the HaF reward model only shows a significant advantage over the baseline model primarily in the middle stages of training. This is likely because both models have largely achieved harmless responses on the test set, resulting in minimal differentiation between the two reward models.
@@ 5 Related Work
Reward model was proposed to modeling human language preferences (model that outputs preference values based on questions and answers) (Christiano et al., 2017), then the explosive growth of research on reward models (McKinney et al., 2023) and large language models (Wei et al., 2022; Park et al., 2023; Zheng et al., 2023; Yue et al., 2024) emerged after the popularity of ChatGPT.
@@
Although many data-oriented methods have greatly enhanced the performance of reward models, the field of reward model optimization has been rarely explored. Currently, the training of reward models basically follows the process proposed by OpenAI (Christiano et al., 2017). Considering the widespread practical applications of reward models, the attention given to their training paradigms does not match their importance.
@@ Appendix A Experiments Setup
lora rank 64 optimizer AdamW precision bf16
@@
To train the reward model, we use DPO Loss as the policy loss in HAF and set policy ratio $\alpha=0.2$. The learning rate is $1.0\times 10^{-5}$ for Phi-2 and Mistral-lora-baseline, $3.0\times 10^{-5}$ for Mistral-lora-HAF. A single RTX A6000 with 48GB memory is used for training the reward model. The model used for testing is the checkpoint that achieves the highest reward on the validation set.
For PPO training in Section 4.2.2, we utilize two RTX A6000 GPUs for parallel training with a total batch size of 4. The maximum number of new tokens generated is set to 128, and the learning rate is 1e-6. The training is conducted over a maximum of 20,000 episodes. We employ score scaling and score normalization and clip the scores between -3 and 3. All other settings follow the implementation in the TRL library. The model used for testing is the checkpoint that achieves the highest reward on the validation set. The generation config includes $top\_p=0.8$, $temperature=0.5$, $length\_penalty=1.3$, $repetition\_penalty=1.2$, $do\_sample=True$
@@ Appendix F GPT Judgement
Table 8 shows the consumption approximation for getting top-1, top-2 responses and the complete order out of 4/8 responses. We consider that performing a single sorting operation on eight responses with the model may result in a loss of precision. Besides, while binary comparisons exhibit high accuracy, repeated binary comparisons inevitably lead to cumulative errors and erroneous outcomes. Therefore, whether from a cost or accuracy standpoint, it is not a favorable option. In practice, we obtain the top 2 responses by ranking 4 responses with GPT-3.5-turbo at once. For 8 candidate responses, we first evenly divide them into two groups and use GPT to rank the responses of each group, then we rank the two sets of the top 2 responses to get the top 2 responses among 8 candidates.
@@
Phi-2HAF 28.68 52.51 32.69 53.35 37.52 66.21 45.44 74.26 24.52 45.15
Phi-2baseline 15.46 34.64 29.28 49.72 27.83 51.68 43.62 73.92 17.22 37.29
MistralHAF 17.42 31.22 9.94 17.70 16.00 28.81 13.68 27.57 9.50 21.07
Mistralbaseline 10.97 23.87 7.45 17.08 17.99 32.78 12.68 26.36 8.68 17.36